In 8 years time Tess will be 5 times as old as her age 8 years ago.
Answer by thejackal(72)

lets call Tess's age today T
thus, 8 years tess's age was T - 8
and thus 8 from now (8+T) tess will be 5 times her age 8 years ago (5(T-8))
so how old is Tess?
well, we now have a beautiful linear equation with only one unknown
given as 8+T = 5(T - 8)
simplify to get 8+T = 5T - 40 which leads to 8 + 40 = 5T - T or 48 = 4T
i'm sure you can work out T on you own now